Dexamethasone is an anti-inflammatory steroid used to treat various allergic disorders, as well as symptoms arising from neurological disorders, such as multiple sclerosis (MS). For a pediatric patient, a typical initial dose is 5 mg/day per square meter of body surface area (BSA). If the BSA of the patient is 1.2 m2, how many 0.75-mg tablets will be administered in a week?
Answer #1

ANSWER.

* Calculation of number of tablet required for the patient.

      Ordered dose = 5 mg/day/BSA.

        BSA of the patient = 1.2 m2.

      So

        Dose required for the patient with BSA 1.2 m2 per week

                                        = 1.2 * 5 * 7

                                        = 6 * 7

                                        = 42 mg.

    Dose of one tablet = 0.75 mg.

    Number of tablets required for one week = Total dose required / Dose of one tablet.

                                                                      = 42/0.75

                                                                      = 56 tablets.